OpenStudy (magbak):

I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Using complete sentences, explain the limitations on the base of an exponential function. Using complete sentences, explain how the graphs of exponential functions are different from the other functions you’ve studied (e.g., linear, quadratic, cubic, rational, Etc.)

OpenStudy (magbak):

@whpalmer4 @Hero

OpenStudy (anonymous):

Other graphs have achieves roots or points in the graph where it reachs the x axis, zero and goes into negative numbers

OpenStudy (anonymous):

The exponenetial graph, given you haven't made it such that it isn't negative will have an asymptote of y=0 since 2^(-500)=1/2^(500) which is a very small number but these numbers will never quite reach 0, but get very close

OpenStudy (magbak):

This is for which question?

OpenStudy (anonymous):

the second one , you can also read here its a much more thorough and visual explanation of what i was saying about the asymptote http://www.purplemath.com/modules/expofcns.htm

OpenStudy (magbak):

So your answer if for the second question.

OpenStudy (magbak):

Ok I just want the answer because the answer explains the entire thing for me please.

OpenStudy (magbak):

@whpalmer4

OpenStudy (magbak):

@Directrix

OpenStudy (magbak):

@jim_thompson5910

OpenStudy (mertsj):

If you look at your book where you studied exponential functions it will tell you what the base has to be. That will answer the first part.

OpenStudy (mertsj):

Or look online for the definition of an exponential function.

OpenStudy (magbak):

I have no book that is the problem!!!

OpenStudy (mertsj):

An Exponential Function is a function of the form y = ab^x, where both a and b are greater than 0 and b is not equal to 1.

OpenStudy (magbak):

Ok what is the limitations on the base of an exponential function.

OpenStudy (mertsj):

Did you read what I posted?

OpenStudy (magbak):

Yes but I am confused please.

OpenStudy (mertsj):

What is the base in the exponential function I posted?

OpenStudy (magbak):

It is x

OpenStudy (mertsj):

\[3^4\]

OpenStudy (mertsj):

What is the base?

OpenStudy (magbak):

Now I am confused.

OpenStudy (mertsj):

in this exponential expression: \[3^4\] 3 is the base and 4 is the exponent

OpenStudy (mertsj):

in this exponential equation: \[y=6(3^4)\] 3 is the base and 4 is the exponent

OpenStudy (mertsj):

In this exponential equation: \[y=ab^x\] What is the base?

OpenStudy (mertsj):

Ok. You keep disappearing. I am done.